I need help with these questions.

Question 6 2 points Save Answer Suppose the population of Friendly Town on September, 2014 is 45,000. The number of new active cases of gonorrhea for May 1st through September 30th is 34 and the number of active cases on September 30th of gonorrhea is 234. The prevalence rate was 76 per 100,000 population 520 per 100,000 population 596 per 100,000 population 7.6 per 100,000 population 52 per 100,000 population
Answer #1

Q.6

prevalence rate means:

it means a number of cases affected in a given time period. the formula for calculating prevalence rate is

prevalence rate = No. of existing cases in a given time period / No. of population during this period X 100,000 population

where as no of existing cases means= new active case and active case=34+234=268

no of population during this time perod is = 45000

SO, PREVALENCE RATE IS= 268 / 45000 X 100,000

=595.555

~596

So, the prevalence rate is 596 per 100,000 population.